OpenStudy (anonymous):

Hercules Films is deciding on the price of the video release of its film Bride of the Son of Frankenstein. Marketing estimates that at a price of p dollars, it can sell q = 300,000 − 15,000p copies, but each copy costs $4 to make. What price will give the greatest profit?

OpenStudy (anonymous):

So, for a given price, p, we'll sell q units, at a profit of P. q=300000−15000p P=q∗(p−4)=(300000−15000p)(p−4) If you expand P and find the vertex, you'll have the best price. If you can use calculus, take the derivative of P with respect to p and set it equal to 0. Solve for p and you have your answer.